Fermi National Laboratory
 
Operation Summary for February 29, 2012 - March 2, 2012
 

Wednesday February 29

The day shift began with Operations monitoring the Linac, Booster, and Main Injector (MI), and with the NuMI and MiniBooNE experiments taking beam.

During the day shift, Preaccelerator personnel made a ten-minute access into the H- Source at 9:49 AM. Operators loaded protons into the Recycler at 10:17 AM. Operators prepared Neutrino Area enclosure N01 for and access at 11:24 AM. A Booster RF station (BRF2) tripped off at 11:30 AM; it reset. At 11:43 AM, a Linac RF station (LRF1) tripped off; it reset. Operations disabled beam to MiniBooNE at 11:49 AM in preparation for an access later in the day. Operations established beam to the Meson FTBF experiment T-1015 around 11:55 AM. EE Support personnel investigated a possible current ripple in a Switchyard (SY) power supply (V714) at 1:10 PM that was causing problems for the MTest beam; this investigation held off beam to T-1015. A Water Group person reported at 1:29 PM that he had isolated the MI LCW system from CUB. A MI person began a g-2 (g minus 2) study at 1:22 PM. Operations established beam to MiniBooNE at 2:40 PM. Operations reported at 3:14 PM that a MI low-level RF CAMAC crate had stopped responding; the MI Specialist said that this might have been related to a Tevatron crate that was turned off earlier. Operators turned on the Muon Rings for a study at 3:17 PM.

During the evening shift, the MI Specialist reset the MI batch intensity chassis at 4:06 PM, which was holding off the NuMI beam permit. LRF1 tripped off at 4:54 PM; it reset. At 5:18 PM, a Linac Klystron RF station (KRF1) tripped off; it reset. The g-2 study ended at 5:32 PM. Operations reported that the output of LRF5 had began drifting at 6:46 PM. The Muon Rings study ended at 7:07 PM. Operations disabled beam to T-1015 at 10 PM.

Thursday March 1

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 12:29 AM, LRF4 tripped off; it reset. At 4:53 AM, the MINERvA experimenter reported that several devices had been reading zero; Operators investigated and reported that beam had been consistent all shift despite the readings and the problem was probably due to a MI-3 CAMAC link. Operations notified Controls personnel about the link problem. At 7 AM, Neutron Therapy Facility (NTF) personnel prepared to treat three patients later in the morning.

During the day shift, EE Support personnel reported at 8:10 AM that they would be changing the taps for a SY power supply (V714) in an attempt to fix the ripple in the MTest beam. Operations reported at 8:22 AM that a scheduled Feeder 24 power outage for Muon Rings maintenance work. EE Support personnel completed their work on V714 at 8:52 AM. A control person rebooted the MI-3 node at 9:55 AM. Operations established beam to T-1015 at 10:37 AM. At 12:01 PM, three Linac quadrupole magnet power supplies tripped off; they all reset. At 12:10 PM, LRF1 tripped off; it reset. At 3:19 PM, BRF3, BRF11, and BRF19 tripped off; they reset.

During the evening shift, Operators reported that FESS personnel restored the Feeder 24 power at 4:26 PM; Operators restored the Muon Rings CAMAC link and turned on vacuum pumps. At 4:57 PM, KRF3 tripped off; it reset. At 5:56 PM, LRF1 tripped off; it reset. KRF5 tripped off at 7:25 PM; it reset. At 8:06 PM, LRF4 tripped off; it reset. At 9:08 PM, LRF3 tripped off for the first of four times in about 35 minutes; it reset, but Linac personnel came in and adjusted the station. Operations disabled T-1015 beam at 10:01 PM.

Friday March 2

The midnight shift began with Operations sending beam to NuMI and MiniBooNE.

During the midnight shift, at 1:04 AM, LRF3 tripped off for the first of two times in 43 minutes; it reset. At 1:14 AM, Operators responded to a LRF5 alarm and filled the RF side and cavity side with LCW. KRF5 tripped off at 3 AM; it reset. At 6:03 AM, a Booster bulk power supply (B:CEP400) tripped off; it reset. At 7:45 AM, NTF personnel prepared to treat two patients later in the morning.

Operations Coordinator (OpsCo) Plans for Today and the Weekend

The OpsCo agrees with the MI and SY personnel plans to conduct studies. Operations received the Run Conditions for sending beam to the Neutrino Area SeaQuest experiment (E906), but the experiment does not yet have the required Operational Readiness Clearance to receive beam.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 03/02/2012    email Fermilab
Security, Privacy, Legal  |  Use of CookiesFermi National Accelerator Laboratory